### v1

# StableSketcher: Enhancing Diffusion Model for Pixel-based Sketch Generation via Visual Question Answering Feedback

Abstract
Although recent advancements in diffusion models have significantly enriched the quality of generated images, challenges remain in synthesizing pixel-based human-drawn sketches, a representative example of abstract expression. To combat these challenges, we propose StableSketcher, a novel framework that empowers diffusion models to generate hand-drawn sketches with high prompt fidelity. Within this framework, we fine-tune the variational autoencoder to optimize latent decoding, enabling it to better capture the characteristics of sketches. In parallel, we integrate a new reward function for reinforcement learning based on visual question answering, which improves text-image alignment and semantic consistency. Extensive experiments demonstrate that StableSketcher generates sketches with improved stylistic fidelity, achieving better alignment with prompts compared to the Stable Diffusion baseline. Additionally, we introduce SketchDUO, to the best of our knowledge, the first dataset comprising instance-level sketches paired with captions and question-answer pairs, thereby addressing the limitations of existing datasets that rely on image-label pairs. Our code and dataset will be made publicly available upon acceptance.
footnotetext: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This work has been submitted to IEEE Access for possible publication.
1 Introduction
The advent of diffusion models has redefined paradigms in text-to-image synthesis, achieving remarkable photorealism [24]. Despite their success in generating detailed images, existing diffusion models exhibit significant shortcomings in synthesizing abstract art forms like sketches. Sketches, as a concise yet intuitive medium for visual expression, offer a unique method of abstract representation by distilling complex ideas into fundamental visual forms. This simplicity makes sketches an ideal form for generative models to emulate abstract reasoning [34].
Figure 1: An overview of our StableSketcher framework and SketchDUO dataset.
The application of sketches spans diverse domains, including sketch-guided text-to-image generation [32, 39, 28], sketch-guided image editing [20, 35], and image retrieval [14, 25, 3], underscoring their significance in both creative and practical contexts. However, generative models often fail to capture the essence of human-drawn sketches, instead generating hyper-realistic renderings that deviate from the simplicity and abstraction inherent in sketches. Moreover, these models struggle with maintaining prompt fidelity, as illustrated in Figure 1.
To address these challenges, we propose StableSketcher, a framework that enhances the generative performance of Stable Diffusion [23] for abstract, human-drawn sketches. We fine-tune the variational autoencoder (VAE) of Stable Diffusion to optimize latent representations, ensuring stylistic coherence in generated outputs. Additionally, we define a novel reward function based on visual question-answering (VQA) feedback, integrating it into a reinforcement learning (RL) algorithm to improve the prompt fidelity of the generated sketches. Qualitative and quantitative evaluations, along with user studies, demonstrate that our framework outperforms the Stable Diffusion baseline in generating abstract sketches with improved prompt fidelity.
Along with the outlined issues, the development of robust sketch generation models has been hindered by the inherent limitations of existing sketch datasets [4, 5, 27, 21]. These datasets lack the semantic depth required for generative tasks, making them fit for sketch classification, but insufficient for text-to-image tasks. Furthermore, existing resources lack the fine-grained, instance-centric sketch–caption pairs required for sketch generation. Existing caption datasets [2, 22] describe relationships across multiple objects in a scene rather than the instance itself.
To combat these limitations, we propose SketchDUO, a comprehensive dataset containing 35.8K instance-level sketches paired with fine-grained textual captions and 54.3K question-answer (QA) pairs, offering rich semantic detail for modeling single-object sketches. SketchDUO includes both positive examples, reflecting the desired sketch style, and negative examples that capture common misrepresentations observed in Stable Diffusion outputs, such as sketches with excessive detail or shading. By incorporating contrastive examples, SketchDUO enhances the model’s understanding of desired and undesired styles, enabling it to generate sketches that better align with the intended style and fidelity.
We summarize the contributions below:
  • We propose StableSketcher, a pixel-based sketch generation framework that adapts Stable Diffusion to generate abstract, human-drawn, instance-level sketches with improved stylistic and prompt fidelity.
  • We introduce a new VQA-based RL reward function to improve semantic alignment with textual prompts. Furthermore, we propose a loss function for optimizing the VAE of Stable Diffusion, enhancing reconstruction quality.
  • We present SketchDUO, a dataset comprising instance-level sketches paired with fine-grained textual captions and QA pairs, highlighting desired and undesired styles through positive and negative examples to reflect a contrastive approach.
2 Literature Review
In this section, we first review the literature on sketch generation with diffusion models, followed by a discussion on sketch datasets and the application of reinforcement learning in diffusion models.
2.1 Sketch Generation with Diffusion Models
Diffusion models are typically trained on large datasets of photorealistic images, resulting in a bias towards generating realistic, highly detailed outputs. This training bias limits their ability to generate abstract representations, such as sketches [26, 37, 15]. Furthermore, conventional diffusion models often lack fine-grained control over the structural and abstract elements in sketches, making it difficult to achieve the desired level of simplicity and abstraction [15, 33].
At the same time, the majority of sketch generation research has focused on vector-based and stroke-based approaches [5, 31], which capture sketches at the granularity of individual strokes. While such methods offer computational efficiency, they struggle to handle more complex and detailed sketches, often failing to capture the essence of human-drawn art. More recent work has pivoted to pixel-based generation using diffusion models. For instance,  [9] proposed a scale-adaptive diffusion model for sketch generation, employing a multi-step sampling technique to enhance the quality of the generated sketches. However, their method is constrained by the use of image-caption pairs, limiting its ability to effectively capture the desired style and characteristics of sketches.
Table 1: Comparison of sketch datasets. Prior sets target recognition; SketchDUO uniquely provides both captions and QA pairs.
Dataset # Classes # Sketches/ Class Total # Sketches Caption QA
TU-Berlin [4] 250 80 20K
Sketchy [27] 125 avg. 600 75K
QuickDraw [5] 345 avg. 144K $\sim$50M
SEVA [21] 128 avg. 703 90K
SketchDUO (ours) 30 avg. 1.2k 35.8K
2.2 Sketch Datasets
As research in generative models progresses, a growing variety of sketch datasets has emerged to support advancements in sketch-related studies. Table 1 compares several existing sketch datasets. QuickDraw [5] is one of the largest datasets for sketch classification, but the absence of annotations and low-quality sketches limit its applicability to generative tasks. TU-Berlin [4] provides more complex sketches, but it also lacks a description of instances. Sketchy [27] pairs sketches with images, but the complexity of the sketches and the reliance on image-label pairs limit its usability for sketch generation. More recent efforts, such as SEVA [21], use CLIPasso [31] to generate stroke-based sketches, yet these datasets still rely on photo-sketch pairs and lack the fine-grained, instance-level captions required for high-quality sketch generation.
2.3 Reinforcement Learning in Diffusion Models
The integration of reinforcement learning (RL) with generative models has garnered significant attention in recent years, particularly in the context of text-to-image generation. Reward-weighted maximum-likelihood estimation (RWR) has been widely used to align generated images with textual prompts [16]. However, methods that rely heavily on rewards can suffer from learning instability. To address this issue, denoising diffusion policy optimization (DDPO) [1] introduced a policy gradient-based reinforcement learning method that optimizes diffusion models directly, enabling users to define custom reward functions tailored to specific generative tasks. In DDPO, BERTScore [40] is used as a text-image alignment reward, but this approach is limited by the methodology that BERTScore relies on image captioning models, which hinders stability when representing abstract forms like sketches.
We leverage the TIFA score [10], which employs a question-answer set-based evaluation method to facilitate a more fine-grained assessment of text-image alignment. Expanding upon the TIFA score, we devise a novel VQA-based reward function, which is incorporated as feedback during training.
3 SketchDUO
SketchDUO contains 35,851 instance-level sketch images paired with textual captions and 54,370 question-answer (QA) sets. By offering both captions and QA pairs, SketchDUO provides rich descriptions for individual objects, effectively addressing the limitations of existing datasets. The dataset adopts a contrastive approach, featuring positive examples that capture the desired sketch style, and negative examples that highlight common misrepresentations in Stable Diffusion outputs, such as images with excessive detail, over-shading, or sketches that resemble photographs of pencil drawings rather than true hand-drawn representations.
  1. 1.
    Fashion Items: Hat, Shoe, T-shirt, Umbrella.
  2. 2.
    Animals: Butterfly, Cat, Cow, Dog, Elephant, Fish, Horse, Rabbit.
  3. 3.
    Nature & Environment: Flower, Leaf, Moon, Sun, Tree.
  4. 4.
    Fictional Characters & Symbols: Angel, Mermaid, Snowman, Teddy Bear
  5. 5.
    Fruits & Food: Apple, Banana, Cake, Pineapple, Strawberry.
  6. 6.
    Household Items: Alarm Clock, Bicycle, House, Mug.
Figure 2 visualizes the category and class distributions, including proportions across categories, class allocations, and sample distributions in the positive and negative datasets.
3.1 Definition of a Sketch
We define a sketch as a simple, human-drawn representation of a single instance. The sketch is characterized by a black line drawing on a white background with no texture, capturing the essence of the object with minimal complexity. The representation must be instance-level, focusing on a single object without excessive details or unnecessary elements.
Figure 2: (a) Proportional distribution of the six categories in SketchDUO, shown as the number of categories and their respective percentages. (b) Number of data samples within each category in SketchDUO, displayed as counts and their respective percentages. (c) Class-level percentage distribution in the positive dataset. (d) Class-level percentage distribution in the negative dataset.
Table 2: Examples from the SketchDUO dataset, showcasing both positive and negative sketches along with their corresponding captions and question–answer pairs.
Category Positive File Name fish_49.png
[Uncaptioned image] Caption A simple drawing of a fish with three curved lines on its body and a round eye on a white background.
Instance Q&A
$Q_{1}$: What animal is in the picture?
$A_{1}$: Fish
$Q_{2}$: How many lines are on the fish?
$A_{2}$: 3
Sketch Q&A
$Q_{1}$: Is the background white?
$A_{1}$: Yes
$Q_{2}$: Is this a simple or a complex drawing?
$A_{2}$: Simple
Category Negative File Name fish_2.png
[Uncaptioned image] Caption A detailed drawing of a blue and red fish with orange accents on a beige background featuring a lot of shading.
Instance Q&A
$Q_{1}$: What color is the fish?
$A_{1}$: Blue and red
$Q_{2}$: Are there orange accents on the fish?
$A_{2}$: Yes
Sketch Q&A
$Q_{1}$: Is there a detailed drawing?
$A_{1}$: Yes
$Q_{2}$: Is there a lot of or a little shading?
$A_{2}$: A lot of
Table 2 presents an example of images, captions, and QA sets from SketchDUO. To construct SketchDUO, we selected 30 common classes shared between the QuickDraw [5] and TU-Berlin [4] datasets. The selection of classes was designed to achieve a balanced representation of diverse objects, ensuring broad thematic coverage across the dataset.
SketchDUO comprises 30 classes, distributed across six broad categories to ensure balanced representation and thematic diversity. Below, we outline the six main categories, their corresponding classes, and their relative distributions:
3.2 Sketch Image Collection
Figure 3: Background-preserving sketch augmentations: original; rotation by $15^{\circ}$ with white padding; rotation by $15^{\circ}$ with white padding; Gaussian blur (weak, $k=3$, $\sigma=0.8$); Gaussian noise applied to strokes only (weak, $\sigma=8$); Gaussian blur (strong, $k=5$, $\sigma=1.6$); Gaussian noise applied to strokes only (strong, $\sigma=16$); and line thickening obtained by binarization followed by morphological dilation with a $3\times 3$ elliptical kernel for one iteration.
We curate a corpus of $24,000$ positive and $11,851$ negative sketch images. The positive portion is derived from $3,000$ human-drawn sketches spanning 30 classes, with one hundred sketches per class. The negative portion contains $1,693$ images generated with Stable Diffusion v2.1 and selected for off style traits such as intricate line work, the presence of color, nonwhite backgrounds, and heavy shading. To every image we apply the same seven background-preserving augmentations. For negative samples, which contain no strokes, we omit the line thickening augmentation. These include rotations of plus or minus fifteen degrees with white padding; Gaussian blur with a weak setting where $k$ equals three and sigma equals $0.8$, and a strong setting where $k$ equals five and $\sigma$ equals 1.6; Gaussian noise applied to strokes only with a weak setting where $\sigma$ equals eight and a strong setting where $\sigma$ equals sixteen; and line thickening achieved through morphological dilation, as detailed in Figure 3. The resulting collection contains $35,851$ images in total, comprising $24,000$ positive and $11,851$ negative samples. A comprehensive analysis of the distributions at the category and class levels for both subsets, including relative proportions and sample counts, is presented in Figure 2.
3.2.1 Positive Sketch Data Construction
To construct the positive samples, we collect 3,000 hand-drawn sketch images, with 100 instances per class. Each image was created by human participants following a standardized protocol to ensure stylistic and semantic consistency across the dataset. Participants were provided with reference examples that described the desired visual characteristics of positive sketches. All sketches were drawn using a tablet and Microsoft Paint (mspaint), the default drawing tool on Windows, to maintain uniformity in drawing tools and procedures. The images were saved in PNG format with a fixed resolution of $\times512$ pixels. The drawing style was carefully controlled to reflect the core properties of the dataset. Each sketch was required to depict a single object instance in a minimal style using thick black lines (#000000) on a pure white background (#FFFFFF). Drawings were intended to be completed within 40 to 60 seconds to encourage simplicity and consistency in visual abstraction. Captions for each image were initially generated using GPT-4o, and subsequently refined through human correction to improve alignment with the visual content. These finalized captions were then passed to a question generation module adapted from the TIFA framework [10], which leverages LLaMA 2 [30] for generating diverse questions and UnifiedQA [12] for answer validation. This process yielded structured QA sets for each image, ensuring that every positive sample in the dataset is paired with high-quality captions and reliable question–answer triplets.
3.2.2 Negative Sketch Data Construction
To build a high-quality negative sample set for contrastive learning, we constructed 1,693 sketch–caption–QA triplets. All images were generated using Stable Diffusion v2.1, and captions were produced via GPT-4o and subsequently refined through human post-editing. QA pairs were generated using the TIFA framework [10], which combines LLaMA 2 [30] for question generation and UnifiedQA [12] for answer validation. Negative samples were generated through a systematic procedure focused on producing images that diverge from the target sketch style. The prompts were explicitly designed to generate characteristics inconsistent with the black-on-white sketch aesthetic, including detailed line drawings, colored elements, textured or non-white backgrounds, and heavy shading. Following generation, all candidate images underwent human filtering to ensure inclusion criteria and semantic fidelity.
3.3 Sketch-Caption Pair
To construct high-quality sketch–caption pairs, we initially experimented with several state-of-the-art image captioning models. Although models like BLIP-2 Flan T5-xl [18] and mPLUG [17] showed strong performance on general datasets, they often produced generic, incorrect, or overly simplistic captions for sketches. For example, captions such as “a black drawing of a cat” or nonsensical outputs highlighted their inability to capture fine-grained sketch details.
Prompt engineering strategies such as “include three characteristics” and “focus on the eyes” were explored, but they failed to meaningfully improve caption quality. This limitation stems from both the models’ smaller size and their training on object-centric datasets like COCO Captions, which emphasize inter-object relationships over instance-specific descriptions.
To overcome this limitation, larger vision-language models were adopted, namely mPLUG-Owl3 [36] and GPT-4o [11]. These models demonstrated significantly improved ability to generate rich, instance-specific, and semantically faithful captions. For instance, GPT-4o could describe “a playful drawing of a fish with stripes, a small round eye, and a triangular tail fin,” better aligning with the level of abstraction and detail present in our sketches.
All model-generated captions were then refined through a rigorous human correction process to ensure accuracy, especially in counting and spatial relationships [10].
3.4 Sketch-QA Triplet
QA sets are generated using the question generation module from the TIFA framework [10], which combines LLaMA 2 [30] for question generation and UnifiedQA [12] for validating the generated questions. The dataset comprises both positive and negative triplets, with each triplet consisting of a sketch, a corresponding question, and its answer. The positive dataset contains 37,412 QA pairs, while the negative dataset includes 16,958 QA pairs, resulting in a total of 54,370 Sketch-QA Triplets. These triplets are crafted to provide rich semantic detail and understanding of single-object sketches.
4 StableSketcher
Figure 4: Overall architecture of StableSketcher. The input prompt ($T$) is fed into the diffusion model through the CLIP text encoder, where the VAE is fine-tuned using $\mathcal{L}_{\text{recon}}$ and $\mathcal{L}_{\text{LPIPS}}$. Once the diffusion model generates the image ($\hat{I}$), it is passed to the fine-tuned mPLUG-large, along with the question from SketchDUO, to generate the corresponding answer. The VQA-based reward $\mathcal{R}_{\text{VQA}}$, calculated using the TIFAScore, is then used as a feedback signal in the reinforcement learning process.
4.1 Preliminaries
Diffusion Models
Diffusion models are generative models that synthesize data by reversing a gradual noising process [8]. Starting from a clean sample $x_{0}$, Gaussian noise is incrementally added through the forward process $q(x_{t}|x_{t-1})$, until pure noise is reached at step $T$. The model then learns a reverse denoising process $p_{\theta}(x_{t-1}|x_{t})$ to reconstruct data from noise. A common training objective is the noise prediction (score-matching) loss:
$\mathcal{L}_{DM}(\theta)=\mathbb{E}_{x_{0},\epsilon,t}\Big[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\Big],$ (1)
where $\epsilon_{\theta}$ predicts the noise $\epsilon$ added at step $t$. By chaining this reverse process, diffusion models can sample high-quality and diverse outputs from pure noise.
Latent Diffusion Models (LDMs)
Stable Diffusion [23] adapts this framework into a latent space for efficiency. Instead of operating directly in pixel space, an image $x_{0}$ is encoded into a latent representation $z$ via a variational autoencoder (VAE), and the diffusion process is carried out in this lower-dimensional space. The training objective then becomes:
$\mathcal{L}_{LDM}(\theta)=\mathbb{E}_{z,\epsilon,t,c}\Big[\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|^{2}\Big],$ (2)
where $c$ denotes conditioning information such as a textual prompt. By performing diffusion in latent space, LDMs significantly reduce computational cost while maintaining the ability to generate semantically aligned images conditioned on text.
Denoising Diffusion Policy Optimization (DDPO)
While diffusion models generate visually realistic images, they may not align closely with input conditions (e.g., textual prompts) or task-specific objectives. Denoising Diffusion Policy Optimization (DDPO)  [1] addresses this limitation by framing the denoising process as a Markov decision process (MDP) [29], where each denoising step is treated as an action. At timestep $t$, the model predicts a denoised sample $x_{t-1}$ and receives a reward $r(x_{t-1},x_{0},y)$ measuring alignment with the conditioning input $y$. The optimization objective is to maximize the expected cumulative reward:
$\max_{\theta}\;J(\theta)=\mathbb{E}_{x_{0},y}\Bigg[\sum_{t=1}^{T}r\big(x_{t-1},x_{0},y\big)\Bigg].$ (3)
This is optimized using a policy-gradient method adapted to diffusion models. The policy is defined as $\pi_{\theta}(x_{t-1}|x_{t},t,y)$, and the gradient is estimated in REINFORCE style:
$\nabla_{\theta}J(\theta)\approx\mathbb{E}\Bigg[\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(x_{t-1}|x_{t},t,y)\,R_{t}\Bigg],$ (4)
where $R_{t}$ denotes the cumulative return. This framework enables diffusion models to go beyond maximum likelihood training by directly incorporating task-specific feedback, such as prompt fidelity, stylistic constraints, or user-defined rewards. In our work, we employ DDPO with a VQA-based reward function to explicitly improve the semantic alignment of generated sketches with their textual prompts.
Building upon these foundations, we propose StableSketcher, a training framework that adapts Stable Diffusion for sketch generation while incorporating DDPO with VQA-based feedback and evaluation. As illustrated in Figure 4, StableSketcher accepts as inputs a textual prompt $\mathbf{T}$ and a human-drawn sketch $\mathbf{I}$ from the SketchDUO dataset. First, the VAE is fine-tuned on sketch images to improve reconstruction fidelity and stability. Next, the Stable Diffusion UNet is trained with the adapted VAE to synthesize human-drawn–style sketches conditioned on $\mathbf{T}$. Finally, Stable Diffusion is fine-tuned via DDPO using our VQA-based reward, which parses elements from $\mathbf{T}$ and evaluates them individually to strengthen prompt fidelity across samples. The training procedure comprises two main stages executed sequentially: (i) Stable Diffusion training and (ii) VQA-driven DDPO feedback, detailed in the following subsections.
Figure 5: Effect of KL weight on VAE reconstruction quality. In the reconstructed images, $\alpha$ denotes the coefficient in the VAE loss, $\mathrm{L_{recon}}+\alpha\cdot\mathrm{L_{KL}}$.
4.2 Training Stable Diffusion
VAE Fine-tuning for Sketch Reconstruction
The Autoencoder KL [13], used as the frozen VAE in Stable Diffusion, has a loss function composed of two main components. First, the reconstruction loss $\mathcal{L}_{\text{recon}}$ measures how well the input data $x$ has been reconstructed via mean squared error (MSE). This can be expressed as:
$\mathcal{L}_{\text{recon}}=\|x-\hat{x}\|^{2}$ (5)
Second, the Kullback–Leibler Divergence (KL) loss $\mathcal{L}_{\text{KL}}$ evaluates how close the distribution sampled from the latent space is to a normal distribution $\mathcal{N}(0,I)$. A weighting factor $\beta$ is often applied to balance the reconstruction and KL terms:
$\mathcal{L}_{\text{AutoencoderKL}}$ $=\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{KL}}$ (6)
$=\|x-\hat{x}\|_{2}^{2}+D_{\text{KL}}(q(z|x)\|p(z))$ (7)
Using a large KL term over-regularizes the approximate posterior $q_{\phi}(z\mid x)$ toward the standard normal prior, reducing the mutual information $I(x;z)$ and causing posterior collapse, which leads to poor or even failed reconstructions. As shown in Figure 5, collapse occurs when the KL weight is large, while reconstruction becomes feasible again when the weight is reduced to very small values. Conversely, relying solely on pixel-wise reconstruction loss, $L_{recon}$ can result in instability in the loss values, leading to unstable training. In particular, sketch data relies heavily on local and perceptual features such as contours and line thickness, which are difficult to capture with pixel-wise errors alone. Losses like MSE or KL do not adequately reflect these perceptual aspects.
To address this issue, we leverage learned perceptual image patch similarity (LPIPS) [38] as a loss function to better capture the characteristics of sketches. LPIPS measures perceptual similarity based on multi-layer CNN feature maps, capturing not just pixel-level differences but also human-perceived properties such as line sharpness, shape consistency, and visual coherence. This makes it especially suitable for sketch images, where abstraction and contour fidelity are more critical than photorealistic detail. LPIPS is defined as:
$\mathcal{L}_{\text{LPIPS}}=\sum_{l}w_{l}\cdot\|\phi_{l}(x)-\phi_{l}(\hat{x})\|^{2},$ (8)
where $\phi_{l}(\cdot)$ denotes the feature map from the $l$-th layer.
Therefore, our final training VAE loss combines MSE with LPIPS to achieve both stable training and sketch-specific reconstruction quality:
$\mathcal{L}_{\text{VAE}}$ $=\|x-\hat{x}\|^{2}+10^{-1}\cdot\mathcal{L}_{\text{LPIPS}}.$ (9)
UNet Fine-tuning for Text-Aligned Sketch Generation
We perform UNet fine-tuning on Stable Diffusion using sketch–caption pairs from the SketchDUO dataset to adapt the model for generating human-drawn style sketch images. As illustrated in Figure 4, the frozen VAE is replaced with our enhanced VAE to better capture sketch-specific representations. Text prompts are incorporated into the UNet through a cross-attention mechanism, enabling the model to effectively align the denoising process with the given prompt. Furthermore, the denoising diffusion probabilistic models (DDPM) [8] scheduler is employed to ensure a stable and consistent diffusion process during training. We follow the original noise prediction objective of Stable Diffusion [23] for UNet fine-tuning.
4.3 VQA-Guided Fine-tuning with DDPO
Design of VQA-Based Reward Function
DDPO [1] originally employed BERTScore [40] with LLaVa [19] to define a reward signal. However, BERTScore has limitations in capturing fine-grained representations, since it computes similarity based on captions generated by vision-language models (VLMs). In this process, the original image is first converted into a caption, which tends to preserve only coarse, overall semantics while discarding fine-grained visual details. As a result, BERTScore evaluates alignment at a global level but fails to verify whether individual elements of the prompt are accurately reflected in the generated image. To address this, we propose a new reward function inspired by TIFAScore [10], which evaluates the prompt fidelity of text-to-image generation by checking whether each individual element of a text prompt is satisfied by the generated image. Formally, TIFAScore is defined as:
$\text{TIFAScore}=\frac{1}{N}\sum_{i=1}^{N}\delta(f(Q_{i},I),A_{i}),$ (10)
where $N$ denotes the number of question and answer (QA) pairs, $Q_{i}$ is a question derived from the prompt, $I$ is the generated image, $f(\cdot)$ is a VQA model, $A_{i}$ is the ground-truth answer, and $\delta(\cdot)$ is the Kronecker delta function.
VQA-Based Reward Function with SketchDUO QA Triplets
Building on this idea, we design a reward function that captures both instance-level fidelity and sketch-style faithfulness using the sketch–QA triplets from SketchDUO:
$\mathcal{R}_{\text{VQA}}=\alpha\cdot\mathcal{R}_{\text{instance}}+(1-\alpha)\cdot\mathcal{R}_{\text{sketch}}$ (11)
For each image, there are $N+M$ QA pairs, consisting of $N$ instance-related questions and $M$ sketch-related questions. The weighting ratio for each component is controlled by $\alpha$, where $0\leq\mathcal{R}_{\text{VQA}}\leq 1$. We set $\alpha=0.5$ in our experiments:
$\mathcal{R}_{\text{instance}}$ $=\frac{1}{N}\sum_{i=1}^{N}\delta(f(Q^{\text{instance}}_{i},I),A^{\text{instance}}_{i}),$ (12)
$\mathcal{R}_{\text{sketch}}$ $=\frac{1}{M}\sum_{j=1}^{M}\delta(f(Q^{\text{sketch}}_{j},I),A^{\text{sketch}}_{j}).$ (13)
For the VQA backbone, we adopt the mPLUG-large model [17], which achieves strong accuracy among SOTA VQA models with competitive inference time [10].
This reward score $\mathcal{R}_{\text{VQA}}$ is used as the feedback signal in the DDPO training loop. At each training step, the sketch generation model produces candidate images based on text prompts, and $\mathcal{R}_{\text{VQA}}$ is computed by evaluating how well each generated image satisfies the paired questions. A higher $\mathcal{R}_{\text{VQA}}$ score indicates that the image successfully satisfies both semantic correctness and sketch-style intent. This reward guides the policy updates by reinforcing image generations that more faithfully reflect prompt semantics and human-like abstraction.
VQA Model Fine-tuning for Accurate Reward Signals
The accuracy of the VQA model directly impacts the reliability of the reward signal and thus the quality of policy updates. High rewards are assigned when generated sketches align with prompt elements, while mismatches yield lower rewards, guiding the policy toward faithful sketch generation. To improve sketch understanding, we fine-tune the VQA model on the SketchDUO QA set, using 80% of the data for training and 20% for evaluation. A comparison of the baseline mPLUG-large and our fine-tuned model is presented in Table 3.
Table 3: Accuracy comparison of the mPLUG-Large VQA model fine-tuned with SketchDUO.
Baseline (mPLUG-L) Dataset Fine-tuning (Epochs)
2 4 6
Accuracy (%) 61.38 Positive 87.38 88.39 88.80
Both 88.05 89.04 89.38
5 Experiments
5.1 Implementation Details
Dataset: SketchDUO
In this study, a dataset consisting of 3K sketch image-caption pairs, was divided into a training set of 1.8K and a test set of 1.2K, with data utilized differentially across learning stages. In the VAE fine-tuning process, 1.8K sketch images were used for training, while the Stable Diffusion UNet learning employed 1.8K sketch image-caption pairs. The DDPO algorithm-based reinforcement learning was conducted using 1.8K sketch images and approximately 29K corresponding QA pairs. In the process of fine-tuning the mPLUG-large VQA model, additional negative samples were integrated. These comprised about 1.7K sketch images in undesired styles, their corresponding captions, and approximately 17K question-and-answer sets.
Baseline
This paper employs Stable Diffusion v1.5 [23] as the baseline model due to its balanced performance in both image quality and text-image alignment compared to other versions.
Evaluation metrics
To evaluate the quality of the generated images, we adopt five metrics that encompass both image quality and text-image alignment. For image quality assessment, we employ Fréchet Inception Distance (FID) [7] and LPIPS [38]. For text-image alignment, we leverage CLIPScore [6], BERTScore [40], and TIFAScore [10].
Table 4: Quantitative evaluation of generated images using FID, CLIPScore, BertScore, and TIFAScore metrics for different configurations of Stable Diffusion models. The “+ Fine-tuning” rows indicate that fine-tuning was applied to the corresponding base model, while “+ VAE fine-tuning” rows represent the additional application of VAE fine-tuning on top of the fine-tuned model.
Method FID $\downarrow$ CLIPScore $\uparrow$ BertScore $\uparrow$ TIFAScore $\uparrow$
Stable Diffusion v1.5 207.59 $\pm$ 22.3 34.00 $\pm$ 2.6 0.89 $\pm$ 0.03 0.59 $\pm$ 0.15
   + UNet fine-tuning 161.94 $\pm$ 20.3 36.05 $\pm$ 2.6 0.89 $\pm$ 0.03 0.68 $\pm$ 0.13
    + VAE fine-tuning 143.68 $\pm$ 16.6 35.48 $\pm$ 2.5 0.88 $\pm$ 0.03 0.68 $\pm$ 0.12
Stable Diffusion v2.1 230.78 $\pm$ 22.7 31.13 $\pm$ 3.4 0.88 $\pm$ 0.03 0.53 $\pm$ 0.15
   + UNet fine-tuning 144.46 $\pm$ 25.7 34.79 $\pm$ 2.7 0.88 $\pm$ 0.03 0.67 $\pm$ 0.13
    + VAE fine-tuning 172.35 $\pm$ 14.5 34.11 $\pm$ 2.8 0.88 $\pm$ 0.03 0.65 $\pm$ 0.13
Figure 6: BERTScore and TIFAScore evaluations for generated images based on the text prompt describing a ”simple drawing of a pineapple with a crown expressed in 3 layers on a white background.
Preliminary experiments
Initially, we evaluated mPLUG-large on the SketchDUO QA test sets, achieving an accuracy of $61.3\%$, as shown in Table 3. Given the insufficient accuracy of this model, we proceeded to fine-tune mPLUG-large on SketchDUO. The fine-tuned model reached an accuracy of $89.3\%$, significantly improving its performance.
Figure 6 demonstrates that TIFAScore is more suitable than BERTScore for evaluating prompt fidelity, as it better captures the alignment between the text prompt and fine-grained elements of the generated image. While BERTScore focuses on overall semantic similarity, TIFAScore evaluates element-level fidelity, ensuring a more accurate assessment of how well the generated images meet the prompt’s specific requirements.
5.2 Quantitative results
Figure 7: Comparison of the progression of the DDPO algorithm with the new reward function. The left side illustrates the changes in generated images as the DDPO algorithm progresses for two sample prompts using the TIFAScore reward and one sample prompt using the BERTScore reward, while the right graph visualizes the reward progression for the respective prompts over the training steps.
Figure 8: Qualitative comparison of images generated by different models based on the input text prompts. (a) Images generated by Stable Diffusion v1.5, baseline model. (b) Images generated by Stable Diffusion v2.1. (c) Outputs from fine-tuning only the UNet component of Stable Diffusion v1.5. (d) Outputs from fine-tuning both the UNet and VAE components of Stable Diffusion v1.5. “Ours” represents the results from our proposed framework, StableSketcher. “Our dataset” displays the ground truth images corresponding to the text prompts. Each example illustrates a representative class from six categories.
Based on Table 4, Stable Diffusion v1.5 demonstrated superior baseline performance compared to v2.1. While both models showed improvements with UNet fine-tuning, v1.5 achieved greater enhancements in reducing FID and increasing TIFAScore. With additional VAE fine-tuning, v1.5 recorded the lowest FID of $143.68$ and the highest TIFAScore of $0.68$, delivering the best overall results. In contrast, VAE fine-tuning had minimal impact on v2.1’s performance. Therefore, Stable Diffusion v1.5 with UNet and VAE fine-tuning, which offers the best performance in text-image alignment and image quality, was selected for use with the DDPO algorithm.
Figure 7 illustrates the training process of the DDPO algorithm with the proposed reward function. For the text prompts $T_{1}$ and $T_{2}$, the generated images progressively aligned better with the prompts as training progressed. The reward for $T_{1}$ increased from $0.77$ to $0.85$, while the reward for $T_{2}$ improved from $0.59$ to $0.82$. For $T_{1}$, the reward initially increased rapidly and then stabilized, with the generated images progressively reflecting the finer details of the prompt. In contrast, $T_{2}$ showed a steady improvement in the reward function throughout training, and the corresponding generated images consistently aligned more closely with the prompt. Meanwhile, for prompt $T_{3}$, the reward remained nearly unchanged throughout the training process, indicating limitations in achieving full prompt fidelity in the generated images. The right graph of Figure 7 visualizes the reward progression over the training steps, demonstrating that the proposed reward function effectively enhances text-image alignment and stabilizes the learning process. The DDPO algorithm consistently generates images with higher prompt fidelity as training progresses, validating the effectiveness of the proposed reward function.
5.3 Qualitative results.
Figure 8 compares image quality across Stable Diffusion variants and our framework, StableSketcher. Stable Diffusion v1.5 and v2.1 show characteristic failures—v1.5 often over-details or drifts from the text, while v2.1 is more abstract and inconsistent. Fine-tuning the UNet improves simplicity and prompt alignment; tuning both UNet and VAE further boosts fidelity but remains unstable for the “white background” and fully accurate instance generation. In contrast, the proposed StableSketcher applies DDPO, an RL-based policy optimization algorithm, to overcome the limitations of baseline models and achieve the best results. The images generated by StableSketcher resemble human-drawn sketches and faithfully reflect the detailed conditions of the text prompts. Additionally, the results from StableSketcher were the most similar to the ground truth images, demonstrating the effectiveness of the proposed framework. Comprehensive qualitative results for all classes, together with the corresponding text prompts, are provided in the Appendix.
5.4 User study
We conducted a ranking-based user study with 46 participants. For each prompt, participants anonymously compared five model outputs and ranked them from 1, indicating the best, to 5, indicating the worst, along three criteria: Sketch Characteristics, Prompt Fidelity, and Human-Drawn. Table X reports mean ranks, where a lower score indicates a stronger preference. Our method achieved the best mean rank on all criteria, with a score of 1.9 for Sketch Characteristics, 1.7 for Prompt Fidelity, and 1.7 for Human-Drawn, resulting in an overall score of 1.7. Compared to the strongest baseline in column d, which achieved 2.3 for Sketch Characteristics, 2.2 for Prompt Fidelity, 2.2 for Human-Drawn, and 2.2 overall, our method improved the mean rank by 17 percent for Sketch Characteristics and by 23 percent for both Prompt Fidelity and Human-Drawn, yielding a 23 percent relative gain in the overall mean rank. These results indicate consistent user preference for our sketches in terms of stylistic abstraction, textual prompt fidelity, and perceived human-likeness.
Table 5: User study results for each model corresponding to the visual samples in Figure 8. Here, (a) denotes images generated by Stable Diffusion v1.5, (b) denotes images generated by Stable Diffusion v2.1, (c) denotes outputs from fine-tuning only the UNet component of Stable Diffusion v1.5, and (d) denotes outputs from fine-tuning both the UNet and VAE components of Stable Diffusion v1.5.
Criterion (Mean Rank ↓) (a) (b) (c) (d) Ours
Sketch Characteristics 3.8 4.1 2.7 2.3 1.9
Prompt Fidelity 4.3 4.1 2.4 2.2 1.7
Human-Drawn 4.0 4.4 2.5 2.2 1.7
Total Average Rank 4.0 4.2 2.5 2.2 1.7
5.5 Ablation on VAE Loss Functions
We evaluated different loss combinations for VAE fine-tuning through both reconstruction and generation tasks. For reconstruction, input sketches were encoded and decoded; for generation, the fine-tuned VAE was integrated into Stable Diffusion to produce sketches from text prompts (e.g., “A black line drawing of a teddy bear with a friendly smile on a white background.”).
Effect of MSE and LPIPS
When MSE was combined with LPIPS, reconstruction quality improved steadily over 15 epochs, as illustrated in Figure 9. Beyond visual gains, this combination also produced consistent reductions in both pixel-wise and perceptual errors: after 15 epochs, the MSE decreased from $0.0008$ to $0.00017$ and the LPIPS from $0.0028$ to $0.0005$. These results show that MSE preserves low-level accuracy, while LPIPS enforces perceptual consistency in contours and line structures, leading to more stable training and improved sketch reconstruction.
Effect of KL Divergence
When training the VAE with a combination of MSE and KL loss, as in the original formulation, the generated outputs gradually collapsed into almost entirely white backgrounds, as shown in Figure 10. This occurs because a large KL weight over-regularizes the latent space, forcing the encoder to map inputs too closely to a standard normal distribution and thereby discarding fine-grained sketch details. As a result, the model fails to reconstruct the black line structures of the sketches, which is consistent with the findings reported in Appendix G of Stable Diffusion [23].
We also examined the effect of varying the KL weight across several orders of magnitude, from $10^{0}$ down to $10^{-6}$. When the weight was set to $10^{-6}$, the KL divergence loss exploded while the reconstruction loss remained low, indicating unstable training. In contrast, setting the weight to $10^{0}$ caused the KL loss to nearly vanish but led to a collapse in reconstruction quality. These results, illustrated in Figure 5, confirm that improper weighting of the KL term severely degrades the VAE’s ability to preserve sketch information.
Other Loss Variants
We additionally considered binary cross-entropy (BCE) loss; however, its constraint that outputs lie in $[0,1]$ is incompatible with the LPIPS objective and our VAE decoder configuration. Optimizing with LPIPS alone yielded thick yet temporally consistent contours, whereas coupling L1 with LPIPS delivered smaller perceptual gains than the MSE–LPIPS pairing and left LPIPS effectively unchanged ($\approx 1\times 10^{-3}$) after 15 epochs. Taken together, the MSE–LPIPS combination offered the most favorable trade-off between pixel-level fidelity and perceptual sketch quality, as summarized in Figure 11.
Figure 9: Reconstruction quality improvement over 15 epochs with our $L_{VAE}=L_{recon}+10^{-1}\cdot L_{LPIPS}$, the combination of MSE loss and LPIPS loss.
Figure 10: Image generation results across epochs 1 to 15 with MSE and KL combination loss.
Figure 11: Qualitative comparison of reconstruction and generation after 15 epochs under different loss compositions. (a)–(e) correspond to $L_{recon}$, $L_{recon}+10^{-6}\cdot L_{KL}$, $L_{recon}+10^{-1}\cdot L_{LPIPS}$, $L_{recon}+10^{-1}\cdot L_{KL}+10^{-1}\cdot L_{LPIPS}$, and $L_{L1}+10^{-1}\cdot L_{LPIPS}$, respectively.
6 Conclusion
We proposed StableSketcher, a framework that extends Stable Diffusion to human-drawn, pixel-based sketches by fine-tuning the latent autoencoder with a reconstruction–perceptual hybrid loss and introducing a reinforcement learning strategy guided by a VQA-based reward function. To support this, we introduced SketchDUO, the first dataset to provide triplets of sketch images, fine-grained captions, and QA pairs, enabling multimodal learning signals tailored to abstract sketch generation. In addition to this unique triplet structure, SketchDUO also incorporates positive examples that reflect desired abstraction and negative examples capturing common errors such as over-shading, excessive detail, or photorealistic bias, offering a contrastive design that enriches model training. Together, StableSketcher and SketchDUO enable fine-grained prompt alignment and better disentanglement of semantic correctness from stylistic faithfulness.
Experimental results confirmed the effectiveness of our approach. StableSketcher achieved the lowest FID (143.68) and the highest TIFAScore (0.68) across all configurations, outperforming Stable Diffusion v1.5 and v2.1 baselines. Compared to the baselines, StableSketcher reduced FID by up to 30.8% (v1.5) and 25.3% (v2.1), while improving TIFAScore by about 15 – 23% and CLIPScore by 4 – 10%, confirming consistent gains across metrics. In contrast, BERTScore showed little difference, suggesting that while it captures the overall semantic impression of an image, it fails to evaluate whether the generated output accurately and precisely reflects the input textual prompt—highlighting the limitation of caption-based metrics for measuring text-to-image generation prompt fidelity. Ablation studies further demonstrated that the proposed VQA-based reward provided more reliable improvements in prompt fidelity, especially for element-level conditions such as object count and background simplicity. Qualitatively, StableSketcher consistently generated sketches with clearer contour abstraction and reduced noise. A user study further validated its superiority, ranking StableSketcher highest in terms of sketch quality, semantic alignment, and perceived human-likeness.
Two limitations remain with respect to data coverage and evaluation. SketchDUO currently comprises 30 categories and 35.8K sketches, which constrains generalization to long-tail objects, multi-object scenes, and stylistic diversity. To address these limitations, we plan to expand the dataset in both breadth and depth by adding new classes, varying style factors (e.g., stroke density and thickness, contour versus hatching, perspective), and enriching annotations beyond captions and QA to include part labels, keypoints, and per-stroke metadata (order, length, curvature).
Acknowledgment
This research was supported by the ‘Regional Innovation System & Education (RISE)’ through the Seoul RISE Center, funded by the Ministry of Education (MOE) and the Seoul Metropolitan Government (2025-RISE-01-007-05); and the Institute of Information & Communications Technology Planning & Evaluation (IITP) under the Artificial Intelligence Convergence Innovation Human Resources Development grant funded by the Korea government (MSIT) (IITP-2025-RS-2023-00254592). Additionally, this work has been carried out with extensive use of the NEURON computing resource supported by the Korea Institute of Science and Technology Information (KISTI). This study involved human participants. All procedures were approved by the Institutional Review Board (IRB) of Dongguk University (Approval No. DUIRB-2025-05-08) and were conducted in accordance with institutional ethical guidelines.
Appendix: Qualitative Evaluation Results
Figure 12: Qualitative comparison of images generated by different models based on the input text prompts. “Ours” denotes results from the proposed StableSketcher. “Our dataset” shows the ground-truth images corresponding to the prompts.
Figure 13: Qualitative comparison based on input text prompts (set 2). “Ours” denotes results from StableSketcher; “Our dataset” shows the corresponding ground-truth images.
Figure 14: Qualitative comparison based on input text prompts (set 3). “Ours” denotes results from StableSketcher; “Our dataset” shows the corresponding ground-truth images.